OpenStudy (anonymous):

Write the first trigonometric function in terms of the second for θ in the given quadrant. sec θ, sin θ; θ in quadrant I

OpenStudy (dumbcow):

sec = 1/cos cos^2 = 1-sin^2 cos = sqrt(1-sin^2) --> sec = 1/sqrt(1-sin^2)